# Sample-Efficient Reinforcement Learning from Human Feedback via Information-Directed Sampling

### v1

@@ Abstract
We study the problem of reinforcement learning from human feedback (RLHF), a critical problem in training large language models, from a theoretical perspective. Our main contribution is the design of novel sample-efficient RLHF algorithms based on information-directed sampling (IDS), an online decision-making principle inspired by information theory. Our algorithms maximize the sum of the value function and a mutual information term that encourages exploration of the unknown environment (which quantifies the information gained about the environment through observed human feedback data). To tackle the challenge of large state spaces and improve sample efficiency, we construct a simplified surrogate environment and introduce a novel distance measure (named the $\ell_{g}$-distance), enabling our IDS-based algorithm to achieve a Bayesian regret upper bound of order $O(H^{\frac{3}{2}}\sqrt{\log(K(\epsilon))T})$, where $H$ is the episode length, $T$ is the number of episode and $K(\epsilon)$ is related to the covering number of the environment. Specializing to the tabular settings, this regret bound is of order $\tilde{O}(H^{2}\sqrt{SAT})$, where $S$ and $A$ are the numbers of states and actions. Finally, we propose an Approximate-IDS algorithm that is computationally more efficient while maintaining nearly the same sample efficiency. The design principle of this approximate algorithm is not only effective in RLHF settings but also applicable to the standard RL framework. Moreover, our work showcases the value of information theory in reinforcement learning and in the training of large language models.
Index Terms:
Reinforcement learning, Reinforcement learning from human feedback, information-directed sampling, large language models
I Introduction
@@
Reinforcement learning from human feedback (RLHF) is a key technique for aligning large language models (LLMs) to human values [1], and has also shown immense potential in many other fields, such as stock prediction, robot training, medical treatments [2]. It can be viewed as an extension of standard reinforcement learning (RL) in the sense that feedback is not given as a numerical reward, but as a one-bit preference over a trajectory pair. Compared to standard RL, this preference-based setting is often more aligned with real-world scenarios, especially for tasks involving human evaluations [3]. However, a key challenge for applying RLHF algorithms is their reliance on extensive human feedback data, which is usually expensive and time-intensive to collect. To address this challenge, recent works on RLHF mainly focus on developing online learning methods that encourage exploration to improve sample efficiency, thereby reducing the amount of human feedback needed [4]. This brings the RLHF problem back to a fundamental question in RL: how to effectively balance the trade-off between exploration and exploitation to improve sample efficiency?
@@
To tackle this trade-off, two major design principles have been introduced. The first approach, Optimism in the Face of Uncertainty (OFU), typically relies on constructing confidence sets that include the true environment with high probability to construct corresponding policies, one example of which is the Upper Confidence Bound (UCB) approach [5, 6]. In this paper, however, we focus on the less explored second approach, Posterior Sampling, which adopts the Bayesian framework and treats the environment as a random variable. One classical posterior sampling algorithm is Thompson Sampling (TS), which has been proved to be sample-efficient and enjoy sublinear Bayesian regret upper bounds in both RL [7] and RLHF settings [8].
@@
Apart from TS, information-directed sampling (IDS) emerges as a novel and principled online decision-making approach. By incorporating a mutual information term into the policy selection procedure, IDS manages to further encourage exploration about the unknown environment, thus tackling the exploration-exploitation tradeoff to a certain extent [9, 10]. Compared with UCB and TS, IDS is more adept at learning complex information-regret structures, and is more flexible and robust to observation noise [11]. In addition, empirical evidence has demonstrated that IDS performs exceptionally well across a range of scenarios, such as sparse linear bandits [12], bandits with graph feedback [13], Markov Decision Processes (MDPs) [9].
@@
    We first introduce a basic IDS-based algorithm for the RLHF setting where the reward is unobservable and only preference feedback is available (see Sec. IV-A). In each episode, it follows the Bayesian posterior sampling paradigm, and solves an optimization problem that maximizes the sum of an expected value term (exploitation) and a mutual information term (exploration). Here, the mutual information quantifies the amount of information about a learning target (e.g., the environment) that can be gained through the trajectrories and preference.
@@
    To tackle the challenge posed by large state spaces, we construct a simplified surrogate environment as the learning target in our algorithm. Using tools from information theory and posterior consistency theory, we prove that our IDS-based algorithm with surrogate environment (Algorithm 1) achieves a Bayesian regret bound of $O(H^{\frac{3}{2}}\sqrt{\log(K(\epsilon))T})$, where $H$ is the episode length, $T$ is the number of episode, and $K(\epsilon)$ is related to the covering number of the environment. We also specialize our algorithm and results to the tabular RLHF, linear RLHF, and contextual dueling bandit settings, and demonstrate the advantages of our algorithm over existing ones.
@@
    Finally, we propose an Approximate-IDS algorithm (Algorithm 2), which is computationally more efficient than Algorithm 1 while maintaining nearly the same sample efficiency. The advantage of this algorithm is that it does not need to construct the surrogate environment. This algorithm selects policies using an alternative optimization objective that can be optimized with standard RL techniques, such as PPO [14]. Furthermore, we note that the design principle of Algorithm 2 is not only effective for preference-based learning but is also applicable to general RL tasks.
@@
Highlights on technical novelty: In the process of constructing the surrogate environment, we introduce a novel distance measure, the $\ell_{g}$-distance, to quantify the discrepancy between two probability measures (see Eqn. (4) in Sec. IV-B). It retains most of the desirable properties of the $\ell_{1}$-distance, and is crucial for the design of the computationally efficient algorithm (Algorithm 2). Moreover, the use of $\ell_{g}$-distance introduces new analytical challenges not present in previous IDS-related works, which are successfully addressed through a variety of refined analytical tools (see Appendices A and B).
@@
Comparisons with related works: First, we note that most existing works on RLHF assume deterministic rewards, whereas our work considers a more general framework where both transitions and rewards are stochastic. Among existing RLHF algorithms, the most relevant to ours is the TS-based algorithm by [8]. Although in the general setting their algorithm’s regret bound is not directly comparable to ours (as theirs depends on the eluder dimension, while ours depends on the covering number), we note that in the tabular setting, our bound is superior if we coarsely substitute the dimension $d$ with $SA$ in their linear setting. When comparing with prior works on standard RL, we note that our regret bound [^1] We say $f(n)=\tilde{O}(g(n))$ if $f(n)=O(g(n)\cdot\text{polylog}(n))$. $\tilde{O}(H^{2}\sqrt{SAT})$ is superior to the regret bound $\tilde{O}(H^{2}\sqrt{S^{2}A^{2}T})$ of the surrogate-IDS algorithm by [9], even though we consider a more challenging RLHF setting where we rely only on human feedback to learn the reward model. Moreover, compared to a prior work on TS for standard RL [7], our analysis method removes a technical assumption that almost all optimal policies visit almost all state action pairs.
II Related Works
@@
Reinforcement Learning from Human Feedback (RLHF): RLHF has emerged as a critical approach in aligning AI systems with human values, especially in complex tasks where human feedback plays a crucial role [15, 16]. The RLHF framework typically involves a three-stage process: supervised fine-tuning (SFT), reward modeling (RM), and reinforcement learning (RL) using algorithms like Proximal Policy Optimization (PPO)[1, 17]. Direct Preference Optimization (DPO) [18] is another approach that directly uses generative models as reward models and trains them using preference data.
@@
The practical success of RLHF has also sparked a variety of theoretical studies. According to the type of preference feedback, these works can be roughly divided into two categories: action preference [19, 20, 21, 22, 23, 24] and trajectory preference [25, 26, 27, 3, 28, 8]. The literature on action preferences is generally referred to as the contextual dueling bandits. In this paper, we focus on the trajectory preference. Most of the existing work in this area follow the OFU principle with the exception of [8, 23] and [23], who investigate a well-known Bayesian method—TS. Note that [8] uses trajectory preferences and can be applied to the general function approximation framework while [23] focuses on contextual dueling bandits. We also use the posterior sampling method, but unlike TS, our method follows the princple of information-directed sampling.
@@
Information-Directed Sampling (IDS): IDS is a design principle for sequential decision-making problems, which balances exploration and exploitation by evaluating the information gain from each action or trajectory. Ref. [10] first introduces the IDS principle in the bandit setting. They decompose the Bayesian regret into a information ratio term and a cumulative information gain term, and bound the regret by tools from information theory. Based on their work, many studies use this method to analyze the regret of the TS algorithm in bandit settings [29, 30, 31, 32, 33, 12, 13].
@@
Recently, [9, 7] study the Bayesian regret of IDS and TS without any prior assumptions for MDP settings. Ref. [7] focuses on analyzing TS in general settings while [9] proposes a regularized-IDS algorithm for tabular and linear settings. Ref. [11] uses the principle of IDS to design a set of algorithms for multi-agent reinforcement learning. They both use the surrogate environment as the learning target to get a sharper bound. However, implementing the surrogate version of the algorithm is a challenge. In this paper, we introduce IDS into RLHF for general MDP settings. We propose an easy-to-implement surrogate algorithm and prove that the regret upper bound has the same order as the original version.
III Preliminaries
III-A Notations
@@
III-B Finite-horizon MDPs
@@ III-C Interaction protocol
In the RLHF setting, the agent cannot directly receive a numerical reward, but only receives a preference signal $o_{t}$ over trajectory pair $(\tau_{0}^{t},\tau_{1}^{t})$, where $o_{t}$ is a Bernoulli random variable with $\mathbb{P}(o_{t}=1|\tau_{0}^{t},\tau_{1}^{t})\triangleq\mathbb{P}(\tau_{1}^{t}\text{ is preferred to }\tau_{0}^{t})$. We assume the preference follows the Bradley-Terry (BT) model [34], which has been widely used in existing works on RLHF. The BT model assumes the probability of humans preferring one choice to the other is proportional to the exponential of the value of cumulative reward:
@@
In the Bayesian setting, we often need to take conditional expectations with regard to $\mathcal{D}_{t}$. For brevity, we follow the standard notation in [9], letting $\mathbb{P}_{t}(\cdot)\triangleq\mathbb{P}(\cdot|\mathcal{D}_{t})$, and $\mathbb{E}_{t}[\cdot]\triangleq\mathbb{E}[\cdot|\mathcal{D}_{t}]$. Finally, let $\mathcal{R}_{t,h}\triangleq(r_{1}^{t,i},...,r_{h}^{t,i})_{i\in\{0,1\}}$ denote the corresponding potential unobserved rewards, where each $r_{h}^{t,i}$ is a random variable satisfying $r_{h}^{t,i}\sim R_{h}(\cdot|s_{h}^{t,i},a_{h}^{t,i})$.
III-D Value function and Bayesian regret
@@
$R_{T}(\mathcal{E},\pi)\triangleq\sum_{t=1}^{T}V_{1,\pi_{\mathcal{E}}^{*}}^{\mathcal{E}}(s_{1}^{t})-V_{1,\pi^{t}}^{\mathcal{E}}(s_{1}^{t}).$ (1)
@@
$BR_{T}(\pi)\triangleq\mathbb{E}_{\mathcal{E}\sim\rho}[R_{T}(\mathcal{E},\pi)].$ (2)
@@
IV The basic IDS Algorithm
This section introduces a basic IDS algorithm for RLHF settings. In Sec. IV-A, we present the generic form of our algorithm with an abstract learning target. Sec. IV-B suggests constructing a discrete surrogate environment as the learning target and then describes an IDS algorithm with the surrogate environment (Algorithm 1). Sec. IV-C provides the Bayesian regret bound for Algorithm 1, while Sec. IV-D specializes this result to tabular RLHF, linear RLHF, and contextual dueling bandits.
IV-A Algorithm description: a generic form
@@
$\pi_{\mathrm{IDS}}^{t}=\arg\max_{\pi\in\Pi}\mathbb{E}_{t}[V_{1,\pi}^{\mathcal{E}}(s_{1})]+\frac{\lambda}{2}\cdot\mathbb{I}_{t}^{\pi}\left(\chi;(\mathcal{H}_{t},\mathcal{R}_{t,H})\right),$ (3)
where $\lambda>0$ is a tunable parameter. Here, $\chi$ is called the learning target, which is a random variable and is usually selected as the whole environment $\mathcal{E}$ when the state space is not too large. However, in Sec. IV-B where we consider large state space cases, we will construct a surrogate environment as the learning target to achieve tighter regret bounds.
The subscript $t$ in $\mathbb{I}_{t}^{\pi}\left(\chi;(\mathcal{H}_{t},\mathcal{R}_{t,H})\right)$ in Eqn. (3) means that the distributions of $\chi$ and $(\mathcal{H}_{t},\mathcal{R}_{t,H})$ are both conditioned on $\mathcal{D}_{t}$, and the superscript $\pi$ means that $(\mathcal{H}_{t},\mathcal{R}_{t,H})$ are obtained by executing the policy $\pi$. Intuitively, a larger value of $\mathbb{I}_{t}^{\pi}\left(\chi;(\mathcal{H}_{t},\mathcal{R}_{t,H})\right)$ indicates that the data obtained at episode $t$ contains more information about the learning target $\chi$. Accordingly, the introduction of mutual information in the policy selection procedure further encourages exploration about the unknown environment, while the expected value term $\mathbb{E}_{t}[V_{1,\pi}^{\mathcal{E}}(s_{1})]$ promotes exploitation. In this way, our algorithm manages to tackle the exploration-exploitation tradeoff to a certain extent.
IV-B Constructing surrogate environments as learning targets

### v3

@@ Abstract
We study the problem of reinforcement learning from human feedback (RLHF), a critical problem in training large language models, from a theoretical perspective. Our main contribution is the design of novel sample-efficient RLHF algorithms based on information-directed sampling (IDS), an online decision-making principle inspired by information theory. Our algorithms maximize the sum of the value function and a mutual information term that encourages exploration of the unknown environment (which quantifies the information gained about the environment through observed human feedback data). To tackle the challenge of large state spaces and improve sample efficiency, we construct a simplified surrogate environment and introduce a novel distance measure (named the $\ell_{g}$-distance), enabling our IDS-based algorithm to achieve a Bayesian regret upper bound of order $O(H^{3/2}\sqrt{\log(K(\epsilon))T})$, where $H$ is the episode length, $T$ is the number of episode and $K(\epsilon)$ is related to the covering number of the environment. Specializing to the tabular settings, this regret bound is of order $\widetilde{O}(H^{2}\sqrt{SAT})$, where $S$ and $A$ are the numbers of states and actions. Finally, we propose an Approximate-IDS algorithm that is computationally more efficient while maintaining nearly the same sample efficiency. The design principle of this approximate algorithm is not only effective in RLHF settings but also applicable to the standard RL framework. Moreover, our work showcases the value of information theory in reinforcement learning and in the training of large language models.
1 Introduction
@@
Reinforcement learning from human feedback (RLHF) is a key technique for aligning large language models (LLMs) to human values Ouyang et al. (2022), and has also shown immense potential in many other fields, such as stock prediction, robot training, medical treatments Zhu et al. (2023). It can be viewed as an extension of standard reinforcement learning (RL) in the sense that feedback is not given as a numerical reward, but as a one-bit preference over a trajectory pair. Compared to standard RL, this preference-based setting is often more aligned with real-world scenarios, especially for tasks involving human evaluations Chen et al. (2022). However, a key challenge for applying RLHF algorithms is their reliance on extensive human feedback data, which is usually expensive and time-intensive to collect. To address this challenge, recent works on RLHF mainly focus on developing online learning methods that encourage exploration to improve sample efficiency, thereby reducing the amount of human feedback needed Xie et al. (2024). This brings the RLHF problem back to a fundamental question in RL: how to effectively balance the trade-off between exploration and exploitation to improve sample efficiency?
@@
To tackle this trade-off, two major design principles have been introduced. The first approach, Optimism in the Face of Uncertainty (OFU), typically relies on constructing confidence sets that include the true environment with high probability to construct corresponding policies, one example of which is the Upper Confidence Bound (UCB) approach Tossou et al. (2019); Ye et al. (2024). In this paper, however, we focus on the less explored second approach, Posterior Sampling, which adopts the Bayesian framework and treats the environment as a random variable. The Bayesian reinforcement learning is a natural framework for studying in-context RL with LLMs (Transformer-based policies). One classical posterior sampling algorithm is Thompson Sampling (TS), which has been proved to be sample-efficient and enjoy sublinear Bayesian regret upper bounds in both RL Moradipari et al. (2023) and RLHF settings Wu and Sun (2023).
@@
Apart from TS, information-directed sampling (IDS) emerges as a novel and principled online decision-making approach. By incorporating a mutual information term into the policy selection procedure, IDS manages to further encourage exploration about the unknown environment, thus tackling the exploration-exploitation tradeoff to a certain extent Hao and Lattimore (2022); Russo and Van Roy (2014). Compared with UCB and TS, IDS is more adept at learning complex information-regret structures, and is more flexible and robust to observation noise Zhang et al. (2024). In addition, empirical evidence has demonstrated that IDS performs exceptionally well across a range of scenarios, such as sparse linear bandits Hao et al. (2021), bandits with graph feedback Hao et al. (2022), Markov Decision Processes (MDPs) Hao and Lattimore (2022).
@@
    We first introduce a basic IDS-based algorithm for the RLHF setting where the reward is unobservable and only preference feedback is available (see Sec. 4.1). In each episode, it follows the Bayesian posterior sampling paradigm, and solves an optimization problem that maximizes the sum of an expected value term (exploitation) and a mutual information term (exploration). Here, the mutual information quantifies the amount of information about a learning target (e.g., the environment) that can be gained through the trajectrories and preference.
@@
    To tackle the challenge posed by large state spaces, we construct a simplified surrogate environment as the learning target in our algorithm. Using tools from information theory and posterior consistency theory, we prove that our IDS-based algorithm with surrogate environment (Algorithm 1) achieves a Bayesian regret bound of $O(H^{3/2}\sqrt{\log(K(\epsilon))T})$, where $H$ is the episode length, $T$ is the number of episode, and $K(\epsilon)$ is related to the covering number of the environment. We also specialize our algorithm and results to the tabular RLHF, linear RLHF, and contextual dueling bandit settings, and demonstrate the advantages of our algorithm over existing ones.
@@
    In the literature on information-directed sampling (IDS), to the best of our knowledge, there is no efficient implementation of IDS that learns a surrogate environment, which hinders its practical application. We propose an Approximate-IDS algorithm (Algorithm 2) that is computationally more efficient than Algorithm 1 while maintaining nearly the same sample efficiency. The advantage of this algorithm is that it does not need to construct the surrogate environment. This algorithm selects policies using an alternative optimization objective that can be optimized with standard RL techniques, such as PPO Schulman et al. (2017). Furthermore, we note that the design principle of Algorithm 2 is not only effective for preference-based learning but is also applicable to general RL tasks.
@@
Highlights on technical novelty: In the process of constructing the surrogate environment, we introduce a novel distance measure, the $\ell_{g}$-distance, to quantify the discrepancy between two probability measures (see Eqn. (4.2) in Sec. 4.2). The newly proposed distance measure facilitates the design of our computationally efficient algorithm (Algorithm 2), while the KL divergence and $\ell_{1}$-distance are unable to achieve this. However, it simultaneously introduces new challenges for theoretical analysis. We overcome this difficulty by investigating the unique properties of the new distance measure, as elaborated in Appendices A and B.
@@
Comparisons with related works: First, we note that most existing works on RLHF assume deterministic rewards, whereas our work considers a more general framework where both transitions and rewards are stochastic. Among existing RLHF algorithms, the most relevant to ours is the TS-based algorithm by Wu and Sun (2023). Although in the general setting their algorithm’s regret bound is not directly comparable to ours (as theirs depends on the eluder dimension, while ours depends on the covering number), we note that in the tabular setting, our bound is superior if we coarsely substitute the dimension $d$ with $SA$ in their linear setting. When comparing with prior works on standard RL, we note that our regret bound [^1] We say $f(n)=\widetilde{O}(g(n))$ if $f(n)=O(g(n)\cdot\text{polylog}(n))$. $\widetilde{O}(H^{2}\sqrt{SAT})$ is superior to the regret bound $\widetilde{O}(H^{2}\sqrt{S^{2}A^{2}T})$ of the surrogate-IDS algorithm by Hao and Lattimore (2022), even though we consider a more challenging RLHF setting where we rely only on human feedback to learn the reward model. Moreover, compared to a prior work on TS for standard RL Moradipari et al. (2023), our analysis method removes a technical assumption that almost all optimal policies visit almost all state action pairs.
2 Related Works
@@
Reinforcement Learning from Human Feedback (RLHF): RLHF has emerged as a critical approach in aligning AI systems with human values, especially in complex tasks where human feedback plays a crucial role Achiam et al. (2023); Touvron et al. (2023). The RLHF framework typically involves a three-stage process: supervised fine-tuning (SFT), reward modeling (RM), and reinforcement learning (RL) using algorithms like Proximal Policy Optimization (PPO)Ouyang et al. (2022); Ziegler et al. (2019). Direct Preference Optimization (DPO) Rafailov et al. (2024) is another approach that directly uses generative models as reward models and trains them using preference data.
@@
The practical success of RLHF has also sparked a variety of theoretical studies. According to the type of preference feedback, these works can be roughly divided into two categories: action preference Fürnkranz et al. (2012); Saha (2021); Ji et al. (2024); Sekhari et al. (2024); Li et al. (2024); Bai et al. (2025) and trajectory preference Busa-Fekete et al. (2014); Xu et al. (2020); Pacchiano et al. (2021); Chen et al. (2022); Taranovic et al. (2022); Wu and Sun (2023). The literature on action preferences is generally referred to as the contextual dueling bandits. In this paper, we focus on the trajectory preference. Most of the existing work in this area follow the OFU principle with the exception of Wu and Sun (2023); Li et al. (2024) and Li et al. (2024), who investigate a well-known Bayesian method—TS. Note that Wu and Sun (2023) uses trajectory preferences and can be applied to the general function approximation framework while Li et al. (2024) focuses on contextual dueling bandits. We also use the posterior sampling method, but unlike TS, our method follows the princple of information-directed sampling.
@@
Information-Directed Sampling (IDS): IDS is a design principle for sequential decision-making problems, which balances exploration and exploitation by evaluating the information gain from each action or trajectory. Ref. Russo and Van Roy (2014) first introduces the IDS principle in the bandit setting. They decompose the Bayesian regret into a information ratio term and a cumulative information gain term, and bound the regret by tools from information theory. Based on their work, many studies use this method to analyze the regret of the TS algorithm in bandit settings Russo and Van Roy (2016); Dong and Van Roy (2018); Bubeck and Sellke (2020); Liu et al. (2018); Kirschner et al. (2021); Hao et al. (2021, 2022).
@@
Recently, Hao and Lattimore (2022); Moradipari et al. (2023) study the Bayesian regret of IDS and TS without any prior assumptions for MDP settings. Ref. Moradipari et al. (2023) focuses on analyzing TS in general settings while Hao and Lattimore (2022) proposes a regularized-IDS algorithm for tabular and linear settings. Ref. Zhang et al. (2024) uses the principle of IDS to design a set of algorithms for multi-agent reinforcement learning. They both use the surrogate environment as the learning target to get a sharper bound. However, implementing the surrogate version of the algorithm is a challenge. In this paper, we introduce IDS into RLHF for general MDP settings. We propose an easy-to-implement surrogate algorithm and prove that the regret upper bound has the same order as the original version.
3 Preliminaries
3.1 Notations
@@
3.2 Finite-horizon MDPs
@@ 3.3 Interaction protocol
In the RLHF setting, the agent cannot directly receive a numerical reward, but only receives a preference signal $o_{t}$ over trajectory pair $(\tau_{0}^{t},\tau_{1}^{t})$, where $o_{t}$ is a Bernoulli random variable with $\mathbb{P}(o_{t}=1|\tau_{0}^{t},\tau_{1}^{t})\triangleq\mathbb{P}(\tau_{1}^{t}\text{ is preferred to }\tau_{0}^{t})$. We assume the preference follows the Bradley-Terry (BT) model Bradley and Terry (1952), which has been widely used in existing works on RLHF. The BT model assumes the probability of humans preferring one choice to the other is proportional to the exponential of the value of cumulative reward:
@@
In the Bayesian setting, we often need to take conditional expectations with regard to $\mathcal{D}_{t}$. For brevity, we follow the standard notation in Hao and Lattimore (2022), letting $\mathbb{P}_{t}(\cdot)\triangleq\mathbb{P}(\cdot|\mathcal{D}_{t})$, and $\mathbb{E}_{t}[\cdot]\triangleq\mathbb{E}[\cdot|\mathcal{D}_{t}]$. The mean environment $\bar{\mathcal{E}}_{t}$ is defined to satisfy $P_{h}^{\bar{\mathcal{E}}_{t}}(\cdot|s,a)=\mathbb{E}_{t}[P_{h}^{\mathcal{E}}(\cdot|s,a)]$ and $R_{h}^{\bar{\mathcal{E}}_{t}}(\cdot|s,a)=\mathbb{E}_{t}[R_{h}^{\mathcal{E}}(\cdot|s,a)]$ for all $s\in\mathcal{S}$ and $a\in\mathcal{A}$. Finally, let $\mathcal{R}_{t,h}\triangleq(r_{1}^{t,i},...,r_{h}^{t,i})_{i\in\{0,1\}}$ denote the corresponding potential unobserved rewards, where each $r_{h}^{t,i}$ is a random variable satisfying $r_{h}^{t,i}\sim R_{h}(\cdot|s_{h}^{t,i},a_{h}^{t,i})$.
3.4 Value function and Bayesian regret
@@
$R_{T}(\mathcal{E},\pi)\triangleq\sum_{t=1}^{T}V_{1,\pi_{\mathcal{E}}^{*}}^{\mathcal{E}}(s_{1}^{t})-V_{1,\pi^{t}}^{\mathcal{E}}(s_{1}^{t}).$ (3.1)
@@
$BR_{T}(\pi)\triangleq\mathbb{E}_{\mathcal{E}\sim\rho}[R_{T}(\mathcal{E},\pi)].$ (3.2)
@@
4 The basic IDS Algorithm
This section introduces a basic IDS algorithm for RLHF settings. In Sec. 4.1, we present the generic form of our algorithm with an abstract learning target. Sec. 4.2 suggests constructing a discrete surrogate environment as the learning target and then describes an IDS algorithm with the surrogate environment (Algorithm 1). Sec. 4.3 provides the Bayesian regret bound for Algorithm 1, while Sec. 4.4 specializes this result to tabular RLHF, linear RLHF, and contextual dueling bandits.
4.1 Algorithm description: a generic form
@@
$\pi_{\mathrm{IDS}}^{t}=\arg\max_{\pi\in\Pi}\mathbb{E}_{t}[V_{1,\pi}^{\mathcal{E}}(s_{1})]+\frac{\lambda}{2}\cdot\mathbb{I}_{t}^{\pi}\left(\chi;(\mathcal{H}_{t},\mathcal{R}_{t,H})\right),$ (4.1)
where $\lambda>0$ is a tunable parameter. Here, $\chi$ is called the learning target, which is a random variable and is usually selected as the whole environment $\mathcal{E}$ when the state space is not too large. However, in Sec. 4.2 where we consider large state space cases, we will construct a surrogate environment as the learning target to achieve tighter regret bounds.
The subscript $t$ in $\mathbb{I}_{t}^{\pi}\left(\chi;(\mathcal{H}_{t},\mathcal{R}_{t,H})\right)$ in Eqn. (4.1) means that the distributions of $\chi$ and $(\mathcal{H}_{t},\mathcal{R}_{t,H})$ are both conditioned on $\mathcal{D}_{t}$, and the superscript $\pi$ means that $(\mathcal{H}_{t},\mathcal{R}_{t,H})$ are obtained by executing the policy $\pi$. Intuitively, a larger value of $\mathbb{I}_{t}^{\pi}\left(\chi;(\mathcal{H}_{t},\mathcal{R}_{t,H})\right)$ indicates that the data obtained at episode $t$ contains more information about the learning target $\chi$. Accordingly, the introduction of mutual information in the policy selection procedure further encourages exploration about the unknown environment, while the expected value term $\mathbb{E}_{t}[V_{1,\pi}^{\mathcal{E}}(s_{1})]$ promotes exploitation. In this way, our algorithm manages to tackle the exploration-exploitation tradeoff to a certain extent.
4.2 Constructing surrogate environments as learning targets